OpenStudy (anonymous):

graph y< x + 3 on a coordinate graph

jigglypuff314 (jigglypuff314):

well first, do you know how to graph y = x + 3 ? :)

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

i have wrestling this weeke for district 10 champs if i fail the rest of my 4 math lessons ill not be able to compete please help me mam its my dream to be a district champ for school

jigglypuff314 (jigglypuff314):

okay, that's alright :) it is the idea that as you plug in numbers for x, you'll get a number for y which you can use to plot points in a (x, y) format. so for example y = x + 3 if you plug in 0 for x then you get y = 0 + 3 right? :) so y = 3 so by pluging in 0 you get 3 therefore a point on the graph would be (0, 3) follow so far?

OpenStudy (anonymous):

so far so good :)

jigglypuff314 (jigglypuff314):

do you understand how to graph a point like (0, 3) ? :)

OpenStudy (anonymous):

but it says graph the inequality

OpenStudy (anonymous):

oh yes i understand how to graph just thesse graphs look complicated

OpenStudy (anonymous):

i only have 4 questions if youd like to help me :)

jigglypuff314 (jigglypuff314):

I'd love to help :) inequality graphs are just shaded versions of the equation graphs :) which was why I was asking if you knew how to graph the equation :P

jigglypuff314 (jigglypuff314):

to get another point on the line y = x + 3 we can plug in 2 for x so y = 2 + 3 = 5 so when x = 2, y = 5 so another point in the format of (x, y) would be (2 , 5) do you understand how I got that? :)

OpenStudy (anonymous):

yea its kinda clciking

jigglypuff314 (jigglypuff314):

good so graph the two points (0, 3) and (2, 5) then draw a straight line through them, that would be your y = x + 3 line :)

jigglypuff314 (jigglypuff314):



OpenStudy (anonymous):

ty ty can i ask you like 3 more question :)

jigglypuff314 (jigglypuff314):

sure, but are you sure you've understood this one? I hadn't even gotten to the inequality part?

OpenStudy (anonymous):

its clciken lil at a time

jigglypuff314 (jigglypuff314):

lol the way inequalities work is that you would either shade above or below the line so if it's like y > something then you would shade above if it's like y < something then you would shade below this is because " > " means "greater than" so if " y " is greater than something, you would shade above same idea with " < " means "less than"

jigglypuff314 (jigglypuff314):

then I would like to ask it if was a < or a ≤

jigglypuff314 (jigglypuff314):